OpenStudy (anonymous):

Simplify. 68.74 - 3.67 × (4.03 - 2.03)^2 58.06 260.28 54.06 67.88

OpenStudy (love_ranaa):

First do 4.03 - 2.03 which equals 2. 2 x 2 =4 Now the equation will be 68.74 - 3.67 × 4

OpenStudy (love_ranaa):

Do 3.67 x 4 which equals 14.68. Now the equation will be 68.74 - 14.68

OpenStudy (anonymous):

@just_one_last_goodbye

OpenStudy (just_one_last_goodbye):

sorry im late

OpenStudy (anonymous):

68.74−3.67(4.03−2.03)2 =54.06

OpenStudy (anonymous):

68.74−3.67(4.03−2.03)2 =68.74−3.67(2^2) =68.74−(3.67)(4) =68.74−14.68 =54.06

OpenStudy (anonymous):

please metal

OpenStudy (just_one_last_goodbye):

@jordanjamesbay did you understand or no?

OpenStudy (anonymous):

ya thx